Help please! How do I solve by elimination:
x-2y = -14 and 3x + 2y = 6

Answer by jennywecas(28)

it would help if you put the equations like this
3x+2y=6
+ x-2y=-14
4x=-8
x=-2
-2-2y=-14
-2y=-12
y=6
the answer is (-2,6)
sorry i had a not so smart moment